OpenStudy (anonymous):

Hey i need help correcting only five questions so i would appreciate it if you stick with me with those five questions

OpenStudy (anonymous):

Hi

OpenStudy (anonymous):

Hi

OpenStudy (anonymous):

Can you help me correct 2 problems

OpenStudy (anonymous):

I mean 5

OpenStudy (anonymous):

To check if i am right

OpenStudy (kuoministers):

depends how hard they are :P

OpenStudy (anonymous):

Lol its algebra 1

OpenStudy (kuoministers):

ok ill try

OpenStudy (anonymous):

Is it good for you O:

OpenStudy (anonymous):

Ok great!

OpenStudy (anonymous):

Let me get the first question

OpenStudy (anonymous):

Solve, using the substitution method. d + e = 2 d - e = 4 (Points : 1) The solution is (3, 1). The solution is (5, 1). There is no solution. There are an infinite number of solutions.

OpenStudy (anonymous):

Ok here it is

OpenStudy (anonymous):

I belive it is 3,-1

OpenStudy (anonymous):

Am i correct

hartnn (hartnn):

yes. correct.

OpenStudy (anonymous):

Boo yah

OpenStudy (anonymous):

i need four more to correct

OpenStudy (anonymous):

Can you guys stay with me please!

OpenStudy (anonymous):

Solve the given system, using the substitution method. y = 3x - 7 6x - 2y = 12 (Points : 1) (14, 12) (12, 14) There is no solution. There are an infinite number of solutions.

hartnn (hartnn):

you keep on posting questions...i'll keep on returning

OpenStudy (anonymous):

Yay!

OpenStudy (anonymous):

Heres the second one!

hartnn (hartnn):

what u got as answer ?

OpenStudy (anonymous):

iS IT NO SOLUTION???

OpenStudy (anonymous):

Caps

OpenStudy (kuoministers):

yea

hartnn (hartnn):

you are \(\huge \color{red}{\checkmark}\)

OpenStudy (anonymous):

Yay 3 more guys you guys been super

OpenStudy (anonymous):

Ok here one!

OpenStudy (anonymous):

I had to print screen it because it didnt let me copy and paste XD

hartnn (hartnn):

whats your answer ?

OpenStudy (anonymous):

4,-1/2

hartnn (hartnn):

you are \(\huge \color{red}{\checkmark}\)

OpenStudy (anonymous):

Am i correct???

hartnn (hartnn):

next.

OpenStudy (kuoministers):

hmm im working it out too slow :P

OpenStudy (anonymous):

Solve, using the substitution method. y - 2x = 8 16 + 4x = 2y (Points : 1) The solution is (24, 0). The solution is (1, 10). There is no solution. There are an infinite number of solutions.

OpenStudy (anonymous):

Ok

hartnn (hartnn):

whats your answer for that ?

OpenStudy (anonymous):

Umm one second

OpenStudy (anonymous):

Is it there no solution

hartnn (hartnn):

incorrect.

OpenStudy (anonymous):

Dang it

OpenStudy (anonymous):

Ok seriously one second dont leave

OpenStudy (anonymous):

Is it 24,0

hartnn (hartnn):

incorrect.

OpenStudy (anonymous):

No wait its infinite amount of solutions!

hartnn (hartnn):

yes. u know how ?

OpenStudy (anonymous):

So i am correct on this problem! :D

OpenStudy (kuoministers):

yes you are but there is a reason why it has no answer or infinite amount of solutions

OpenStudy (anonymous):

Is it because if you do any operation to this problem we will get the same solution witch means infinite amount

OpenStudy (anonymous):

You guys there

OpenStudy (kuoministers):

if the equation has no variables, then it has an infinite number of solutions. like!! 3 = 3 0 = 0 if the equation winds up with no variables, then you are dealing with no solutions. example: 2 = 3 0 = 5 etc.

OpenStudy (anonymous):

So i am correct that it has infinite numbers of solutions right?

hartnn (hartnn):

yes.

OpenStudy (anonymous):

Sweet!

OpenStudy (anonymous):

Ok last one

OpenStudy (anonymous):

Solve, using the substitution method. y = 2x + 6 2x - y = 6 (Points : 1) (14, 12) (12, 30) There is no solution. There are an infinite number of solutions.

OpenStudy (kuoministers):

so what did you get?

OpenStudy (anonymous):

I got.................

hartnn (hartnn):

.............. is incorrect

OpenStudy (anonymous):

14,12

hartnn (hartnn):

incorrect.

OpenStudy (anonymous):

Lol XD

OpenStudy (anonymous):

Awwww

OpenStudy (anonymous):

Ok one second dont leave

OpenStudy (kuoministers):

we wont leave >.<

hartnn (hartnn):

from 1st equation y-2x=6 from 2nd 2x-y =6 is that possible ?

OpenStudy (anonymous):

There is no solution

OpenStudy (anonymous):

Right

hartnn (hartnn):

you are \(\huge \color{red}{\checkmark}\)

OpenStudy (anonymous):

Woot cause i got x = 6

OpenStudy (kuoministers):

yes

OpenStudy (kuoministers):

umm?

OpenStudy (anonymous):

I mean i got x = 0

OpenStudy (anonymous):

Sorry got to excited put the wrong one

hartnn (hartnn):

u can't get x.

hartnn (hartnn):

there is no solution.

OpenStudy (kuoministers):

x cancels itself out.. can i pls see ur working out?

OpenStudy (anonymous):

Yeah thus X = 0

OpenStudy (anonymous):

Thats how my school Teachers XD

hartnn (hartnn):

no, if it cancels, its not 0.....

OpenStudy (anonymous):

Or it could be 0=0

OpenStudy (kuoministers):

because 0 is still an answer right? This has NO solution not even 0

OpenStudy (anonymous):

Ohhh

OpenStudy (anonymous):

Let me draw how i did it one sec

OpenStudy (anonymous):